OpenStudy (anonymous):

Just reviewing for the AP Test. I can't tell whether my brain isn't working or how to solve this: integration of [6/(n^2 +1]dx

OpenStudy (alexwee123):

HEY IS AP CALC BC HARD?

OpenStudy (anonymous):

Some concepts are confusing, but if you actually do the homework, it's not that bad. It's not exactly an easy class though.

OpenStudy (anonymous):

Use the Equation editor to write equation please \[\int\limits_{}^{}\frac{6}{(x^2 +1)}dx\] Is this your question ??

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

do you know a function whose derivative is \[\frac{1}{x^2+1}\] ??

OpenStudy (anonymous):

think "inverse trig"

OpenStudy (anonymous):

GAH! Trig integration again >.< I just derped pretty hard there. Thanks guys =)

OpenStudy (anonymous):

what shivam said it is \[\tan^{-1}(x)\]

OpenStudy (anonymous):

oh no not hard, you should have the derivatives of inverst trig functions memorized more or less

OpenStudy (anonymous):

\[\int\limits\limits\limits_{}^{}\frac{6}{(x^2 +1)}dx= 6\int\limits\limits\limits_{}^{}\frac{dx}{(x^2 +1)}=6 \tan^{-1} x+c\]

OpenStudy (anonymous):

rocking that latex!

OpenStudy (anonymous):

Yeah. Thanks @satellite73 :)